OpenStudy (anonymous):

helllllllpppppp!!!!!!!!!! a cylinder has a radius of 2x+3 and height of 6x+. which polynomial in standard form best describe the total volume of the cylinder?

OpenStudy (anonymous):

V=pi*r^2*h V=pi*(2x+3)^2*6x

OpenStudy (phi):

do you know the formula for the volume of a cylinder ?

OpenStudy (wwe123):

vol of cylinder =2*pi*r*l =2*3.14*(2x+3)*6x =37.68x(2x+3) =

OpenStudy (anonymous):

That is not the correct formula @wwe123

OpenStudy (anonymous):

yea i know the formula i just dont know how to use it

OpenStudy (anonymous):

@phi

OpenStudy (wwe123):

baba ji ka thullu

OpenStudy (wwe123):

baba ji ka thullu @Samuraisam

OpenStudy (phi):

**yea i know the formula i just dont know how to use it *** First write it down. can you do that ?

OpenStudy (anonymous):

done

OpenStudy (phi):

can you post it ?

OpenStudy (anonymous):

V=pi*r^2*h

OpenStudy (phi):

a cylinder has a radius of 2x+3 that means replace r in your formula with (2x+3) height of 6x+ <-- did you leave something out ? whatever the height is, replace h in your formula with the expression can you do that ?

OpenStudy (anonymous):

yes ok i got it thanks btw how to i give medals out i wanna give you one

OpenStudy (phi):

they want the answer in standard form, which means you have to multiply out that mess

OpenStudy (anonymous):

ok well i got 24pix^3+76pix^2+66pix+9pi is that right

OpenStudy (phi):

what is the height ... 6x+ ? or just 6x

OpenStudy (anonymous):

6x+1

OpenStudy (phi):

\[ V = \pi r^2 h \\ V = \pi (2x+3)^2 (6x+1) \\ V = \pi (4x^2 +12x+9)(6x+1) \] now we do \[ V= \pi \left( (4x^2 +12x+9) 6x + (4x^2 +12x+9)\cdot 1\right) \\ V= \pi \left(24x^3 +72x^2 +54x + 4x^2 +12x +9 \right) \] that simplifies to \[ V= 24\pi x^3 + 76 \pi x^2 + 66 \pi x + 9 \pi \] which matches what you got.

OpenStudy (anonymous):

yay (claps whaaooo) thxz

OpenStudy (phi):

yes, that was a bit painful

OpenStudy (anonymous):

i understand lol i have a test full

OpenStudy (anonymous):

@phi